# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ Abstract
We study continuous-time CRRA portfolio choice in diffusion markets with uncertain estimated coefficients. Nature draws a latent parameter from a given distribution and keeps it fixed; the investor cannot observe this parameter and must commit to a parameter-blind policy maximizing an ex-ante objective. We treat the uncertainty distribution as an inference-agnostic sampling input.
We develop a simulation-only two-stage solver. Stage 1 extends Pontryagin-Guided Direct Policy Optimization (PG-DPO) by sampling parameters internally and computing gradients via backpropagation through time. Stage 2 performs an aggregated Pontryagin projection: it aggregates costates across the parameter distribution to enforce a deployable stationarity condition, yielding a structured correction amortized via interactive distillation.
We prove a uniform conditional BPTT-PMP correspondence and a residual-based policy-gap bound with explicit error terms. Experiments on high-dimensional Gaussian drift and factor-driven benchmarks show that projection stabilizes learning and accurately recovers analytic references, while a model-free PPO baseline remains far from the targets.
@@ 1 Introduction
A central problem in quantitative finance is to allocate wealth dynamically across many risky assets in continuous time. In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and closed-form optimal portfolios follow from HJB analysis; see, for example, Merton (1969, 1971). In realistic markets, however, key inputs—especially expected returns—must be estimated from finite samples, often in high dimensions and under regularization/model selection. Empirically, return predictability is unstable and many predictors deliver limited out-of-sample gains (e.g., Goyal and Welch, 2008; Campbell and Thompson, 2008). These considerations motivate separating diffusion risk (Brownian noise conditional on parameters) from decision-time estimation error (uncertainty in estimated coefficients), a theme that also underlies intertemporal hedging and more conservative allocations in models with predictability/learning (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Campbell and Viceira, 2002; Xia, 2001; Maenhout, 2004).
We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, and an external estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$. We treat $q$ as an input object that summarizes decision-time estimation error about $\theta$. Concretely, $q$ may be obtained from resampling-based approximations (e.g., the bootstrap) (e.g., Efron, 1979; Efron and Tibshirani, 1994), from subsample aggregation and model-selection variability (e.g., bagging), or from asymptotic (sandwich/normal) approximations of an estimator’s sampling distribution (van der Vaart, 1998). Our goal is not to revisit inference, but to optimize decisions given this uncertainty description. Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside a simulator; we do not assume a closed-form density, conjugacy, or a particular parametric family. We seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$.
A key modeling choice is that $\theta$ is latent: Nature draws a fixed but unobserved $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed on $[0,T]$. The investor knows $q$ but does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy. We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a filtering/belief process. This is a one-shot decision-time problem: $q$ is treated as fixed over the trading horizon, and we do not model or hedge subsequent updates of the uncertainty description. Belief-state control (optimal learning/hedging of parameter uncertainty) is conceptually important, but it defines a different and typically far more demanding problem than computing a single deployable $\theta$-blind rule from a fixed $q$ (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
The $\theta$-blind deployability constraint also changes what a first-order optimality condition means. Under full information, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition, hence a $\theta$-conditional feedback map that is infeasible when $\theta$ is latent. Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind, and the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the controlled state process (e.g., Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009). In our portfolio setting, $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$, so $q$-aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projection map. This condition is inference-agnostic: it depends on $q$ only through its role as the ex–ante mixing law.
These features place the problem beyond the reach of classical dynamic programming in the targeted high-dimensional regime. Even in moderate dimensions, HJB solvers face the curse of dimensionality (e.g., Bellman, 1961; Kushner and Dupuis, 2001). Although deep PDE (e.g., Raissi et al., 2019; Sirignano and Spiliopoulos, 2018) and BSDE methods (e.g., Han et al., 2018; Beck et al., 2019) reduce grid reliance, nonlinear portfolio HJBs remain numerically delicate. Furthermore, belief augmentation renders the problem infinite-dimensional (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017). Accordingly, we exclude value-function solvers as practical baselines under our $\theta$-blind interface, validating instead on closed-form benchmarks and simulation-only model-free RL (e.g., PPO).
Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (Huh et al., 2025b, a). PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and uses backpropagation through time (BPTT) to compute exact discrete-time gradients of terminal utility. Pathwise sensitivities computed by BPTT coincide with stochastic costates in PMP, reflecting the classical adjoint/backpropagation correspondence (e.g., LeCun, 1988; Yong and Zhou, 1999). In the latent-parameter setting, we sample $\theta\sim q$ inside the simulator (fixing it along each trajectory), while the policy depends only on observable states. To stabilize learning under heterogeneous $\theta$ draws, we develop uncertainty-aware projected P–PGDPO: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated PMP first-order condition, producing a robust deployable $\theta$-blind rule that can be amortized into a fast policy.
Empirically, in high-dimensional static drift experiments, the projected pipeline improves decision-time accuracy over end-to-end learning, stabilizing aligned regimes. In factor-driven markets where correlation induces hedging demand, it recovers analytic references under the $\theta$-blind restriction, whereas a PPO baseline remains far off. Theoretically, we establish a residual-based policy-gap bound: under mild stability, a small aggregated residual implies the policy is close to a locally optimal deployable rule, up to discretization and Monte Carlo errors.
Our contributions are threefold. (i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated PMP first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$. (ii) We extend PG–DPO to this setting by sampling $\theta$ only inside the simulator and using BPTT to compute exact discrete-time gradients and pathwise sensitivities, and we establish a conditional BPTT–PMP correspondence uniform over $\theta$ on compact subsets of $\Theta$. (iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization.
@@
The remainder of the paper is organized as follows. Section 2 formulates the fixed-$q$ ex–ante problem and derives $\theta$-conditional versus $q$-aggregated PMP conditions together with Gaussian decision-time reference models. Section 3 develops PG–DPO and P–PGDPO for latent $\theta$ and proves the policy-gap bound. Sections 4 and 5 report numerical experiments, and the appendix collects technical proofs and implementation details.
@@ 2.1 Model and ex–ante objective in estimated diffusion markets
where $W$ and $W^{Y}$ are Brownian motions (possibly of different dimension) that may be instantaneously correlated. We write the instantaneous covariance and return–factor cross-covariance as
@@
where $\rho$ is defined by $d\langle W,W^{Y}\rangle_{t}=\rho\,dt$. Thus $\Sigma(y,\theta)\in\mathbb{R}^{d\times d}$ and $\Sigma_{SY}(y,\theta)\in\mathbb{R}^{d\times m}$.
@@
We consider the observable market filtration
@@
Approximating the outer expectation in (8) amounts to sampling $\theta\sim q$ inside the simulator (once per trajectory or once per update), running (1)–(2) under that frozen draw, and updating a $\theta$-blind feedback policy to perform well on average over such draws. This is the setting targeted by the simulation-based PG–DPO and P–PGDPO methods developed in Section 3.
@@
Such a law may coincide with a static input (Section 2.3.1), may arise as a model-implied predictive (prior) propagation of an initial uncertainty through factor dynamics without conditioning on new observations (Section 2.3.2), or may be produced by an external filtering/estimation routine that outputs a time-indexed uncertainty description (Appendix A). Our model and algorithms are conditional on whichever $q$ is supplied at the decision time and always freeze that input over the ensuing horizon.
@@
Fix $\theta\in\Theta$ and suppose, for the moment, that $\theta$ were observable to the controller. In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (3). For clarity, we define the control Hamiltonian map as a function of generic “sensitivity” inputs $(\varphi_{x},\varphi_{xx},\varphi_{xy})$:
$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,\varphi_{x},\varphi_{xx},\varphi_{xy}):=x\,\pi^{\top}b(y,\theta)\,\varphi_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,\varphi_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,\varphi_{xy},$ (10)
where $\varphi_{x},\varphi_{xx}\in\mathbb{R}$ and $\varphi_{xy}\in\mathbb{R}^{m}$. In the full-information HJB interpretation, we evaluate (10) at $(\varphi_{x},\varphi_{xx},\varphi_{xy})=(V_{x},V_{xx},V_{xy})$. In the Pontryagin interpretation used below, the same affine map is evaluated at the corresponding $\theta$-conditional adjoint/sensitivity objects (denoted $(p^{\theta},p_{x}^{\theta},p_{y}^{\theta})$ along a trajectory); in smooth Markov regimes these objects coincide with the indicated value derivatives.
@@
To see why ex–ante aggregation enters the first-order condition, take any $\theta$-blind admissible perturbation $h=\{h_{t}\}_{t\in[0,T]}$ that is progressively measurable with respect to the observation filtration $(\mathcal{F}_{t}^{\mathrm{obs}})$ and square-integrable, and define $\pi^{\varepsilon}:=\pi+\varepsilon h$ for small $\varepsilon$. For each fixed $\theta$, the stochastic maximum principle yields the first-variation identity
@@
where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity/adjoint objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market. (Equivalently, (13) evaluates the same affine map $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ defined by (10), with $(V_{x},V_{xx},V_{xy})$ replaced by the corresponding Pontryagin objects along the controlled trajectory.) Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
@@
Consider the fixed-$q$ ex–ante objective (8) over the $\theta$-blind Markov feedback class $\mathcal{A}^{\mathrm{fb}}$. Assume standard smoothness/integrability conditions ensuring validity of first variations within $\mathcal{A}^{\mathrm{fb}}$ and existence of the associated $\theta$-conditional Pontryagin objects. If $\pi^{\star,\mathrm{blind}}$ is a locally optimal interior policy in $\mathcal{A}^{\mathrm{fb}}$, then (15) holds. Moreover, in the portfolio setting (10), the aggregated stationarity is equivalent to the statewise linear system
@@
The conditional first-variation identity (13) is standard for diffusion control under a fixed parameter $\theta$ (e.g. Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009). Taking the outer expectation over $\theta\sim q$ yields (14). Since $h$ is an arbitrary $\theta$-blind admissible perturbation, vanishing of the first variation at an interior optimum implies (15). For the quadratic portfolio Hamiltonian (10), substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ and introducing (16)–(18) yields the linear system (19) and the projected form (20) whenever $A_{t}$ is invertible. ∎

### v3

@@ Abstract
We study continuous-time CRRA portfolio choice in diffusion markets with estimated and hence uncertain coefficients. Nature draws a latent parameter $\theta\sim q$ at time 0 and keeps it fixed; the investor never observes $\theta$ and must commit to a single $\theta$-blind policy maximizing an ex–ante objective, treating $q$ as a decision-time input.
We propose a simulation-only two-stage solver. Stage 1 (DPO) performs BPTT-based stochastic gradient ascent through an Euler simulator while sampling $\theta$ only inside the simulator. Stage 2 (Pontryagin projection) aggregates costate blocks across $\theta\sim q$ and enforces the $q$-aggregated stationarity condition within the deployable class; the resulting correction can be amortized via interactive distillation. We refer to the full Stage 1+Stage 2 pipeline as PG–DPO.
We prove a uniform conditional BPTT–PMP correspondence and a residual-based policy-gap bound with explicit discretization and Monte Carlo error terms. Experiments on high-dimensional Gaussian drift-uncertainty and factor-driven benchmarks show that projection stabilizes learning and accurately recovers analytic decision-time references, while a model-free PPO baseline remains far from the targets.
@@ 1 Introduction
Dynamic portfolio choice is typically studied under the assumption that investment opportunities are known, yielding Merton-type closed-form rules and their intertemporal extensions (Markowitz, 1952; Samuelson, 1969; Merton, 1969, 1971; Karatzas and Shreve, 1998; Campbell and Viceira, 2002). In practice, expected returns and other inputs are estimated from finite samples, are unstable across specifications, and can be highly uncertain at the time a decision is made (Goyal and Welch, 2008; Campbell and Thompson, 2008; Lettau and Van Nieuwerburgh, 2008; Rapach et al., 2010).
A classical response is learning and belief-state control: investors update posteriors over time and hedge against future revisions in investment opportunities (Kandel and Stambaugh, 1996; Barberis, 2000; Xia, 2001; Brandt et al., 2005). While conceptually appealing, belief-state formulations lead to high-dimensional (sometimes infinite-dimensional) state spaces and are difficult to implement in realistic multi-asset settings (Lakner, 1995; Bensoussan and van Schuppen, 1985; Björk et al., 2010; Pham, 2009).
This paper takes a complementary decision-time perspective centered on deployability:
Given an externally supplied description of parameter uncertainty at the decision time, what constitutes an optimal deployable portfolio rule when the true parameter is latent and cannot be conditioned on during trading?
We treat estimation risk as an exogenous decision-time uncertainty object—a fixed law $q(d\theta)$ over market parameters produced by an external pipeline (e.g., bootstrap, subsampling, or asymptotic approximations) (Efron, 1979; Efron and Tibshirani, 1994; Breiman, 1996; van der Vaart, 1998). Nature draws $\theta\sim q$ at time 0 and keeps it fixed over the horizon, but the investor never observes $\theta$ and must commit to a single $\theta$-blind policy.
This information structure changes the relevant first-order condition. Full-information policies such as the Merton rule are defined conditional on the realized parameter and are infeasible when $\theta$ is latent (Merton, 1969, 1971; Campbell and Viceira, 2002). More fundamentally, optimality must be enforced within the deployable class: admissible perturbations are $\theta$-blind, so stationarity requires the $q$-expectation of the Hamiltonian gradient to vanish along the controlled state process. In portfolio problems, this $q$-aggregated Pontryagin condition reduces to a statewise linear system whose solution defines a canonical deployable projected rule. Averaging full-information optima does not, in general, produce such a rule; optimality does not commute with averaging (Pástor, 2000; Avramov, 2002; Cremers, 2002; Brown, 1978).
These observations lead to three contributions. First, we derive the $q$-aggregated Pontryagin first-order condition for the fixed-$q$, $\theta$-blind problem and show how it induces a deployable projection map in portfolio settings (Pontryagin, 2018; Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009). Second, we propose a simulation-only two-stage solver that avoids dynamic programming, which is infeasible in the high-dimensional uncertain regimes of interest (Bellman, 1961; Kushner and Dupuis, 2001; Han et al., 2018; Beck et al., 2019). Stage 1 (DPO) trains a $\theta$-blind policy by stochastic gradient ascent using exact discrete-time gradients computed by backpropagation through time through an Euler simulator, while sampling $\theta$ only inside the simulator. Stage 2 applies a Pontryagin projection by aggregating costate objects across $\theta\sim q$ and projecting the warm policy onto the deployable $q$-aggregated stationarity condition, yielding a structured correction that can be amortized via interactive distillation (Werbos, 1990; Williams and Zipser, 1989; Pardoux and Peng, 1990; Ma and Yong, 1999; Hinton et al., 2015; Jin et al., 2020). We refer to the full Stage 1+Stage 2 pipeline as PG–DPO. Third, we provide a theory that matches this computational design: we prove a conditional BPTT–PMP correspondence that is uniform over compact parameter sets and includes the costate blocks required by the projection step, and we establish a residual-based policy-gap bound with explicit discretization and Monte Carlo error terms. In Gaussian benchmarks, the projected rule recovers familiar shrinkage and horizon-dependent hedging effects connected to classical return-predictability insights (Kandel and Stambaugh, 1996; Campbell and Viceira, 1999; Kim and Omberg, 1996; Lynch, 2001).
Empirically, in high-dimensional settings the Stage 2 projection stabilizes learning and accurately recovers analytic decision-time references, while a model-free PPO baseline remains far from the targets under the same deployability restriction (Schulman et al., 2017; Wang et al., 2020; Jia and Zhou, 2022). Overall, the paper reframes portfolio choice under parameter uncertainty as a problem of commitment under imperfect information at the decision time, rather than explicit belief-state learning.
@@
Section 2 formulates the fixed-$q$ ex–ante problem, contrasts $\theta$-conditional and $q$-aggregated PMP conditions, and presents Gaussian decision-time reference models. Section 3 develops Stage 1 DPO and Stage 2 Pontryagin projection (PG–DPO) for latent $\theta$ and proves the policy-gap bound. Sections 4 and 5 report numerical experiments, and the appendix collects proofs and implementation details.
@@ 2.1 Model and ex–ante objective in estimated diffusion markets
where $W$ and $W^{Y}$ are Brownian motions (possibly of different dimension) that may be instantaneously correlated. For bookkeeping, let $W$ be $d_{W}$-dimensional, $W^{Y}$ be $d_{Y}$-dimensional, and write
$d\langle W,W^{Y}\rangle_{t}=\rho\,dt,\hskip 18.49988pt\rho\in\mathbb{R}^{d_{W}\times d_{Y}}.$ (3)
We write the instantaneous covariance and return–factor cross-covariance as
@@
so that $\Sigma(y,\theta)\in\mathbb{R}^{d\times d}$ and $\Sigma_{SY}(y,\theta)\in\mathbb{R}^{d\times m}$.
@@
All processes are defined on a filtered probability space $(\Omega,\mathcal{F},\mathbb{F}=(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$ satisfying the usual conditions and carrying $(W,W^{Y},\theta)$.
@@
We consider the observable market filtration (a subfiltration of $\mathbb{F}$)
@@
Notation. Throughout Section 2, $\rho$ denotes the Brownian correlation matrix in (3) (and (33)); $T$ denotes the terminal time/horizon; and $K$ denotes the OU mean-reversion matrix in (31). We do not reuse $\rho$, $T$, or $K$ for unrelated quantities elsewhere.
@@
Approximating the outer expectation in (9) amounts to sampling $\theta\sim q$ inside the simulator (once per trajectory or once per update), running (1)–(2) under that frozen draw, and updating a $\theta$-blind feedback policy to perform well on average over such draws. This is the setting targeted by the simulation-based DPO and PG–DPO methods developed in Section 3.
@@
Such a law may coincide with a static input (Section 2.3.1), may arise as a model-implied predictive (prior) propagation of an initial uncertainty through factor dynamics without conditioning on new observations (Section 2.3.2), or may be produced by an external filtering/estimation routine that outputs a time-indexed uncertainty description (Appendix B). Our model and algorithms are conditional on whichever $q$ is supplied at the decision time and always freeze that input over the ensuing horizon.
@@
Fix $\theta\in\Theta$ and suppose, for the moment, that $\theta$ were observable to the controller. In Markovian settings with sufficient smoothness, the $\theta$-conditional value function $V^{\star,\theta}(t,x,y)$ satisfies an HJB equation whose control Hamiltonian (the part depending on $\pi$) can be written explicitly using (4). For clarity, we define the control Hamiltonian map as a function of generic “sensitivity” inputs $(\psi_{x},\psi_{xx},\psi_{xy})$:
$\mathcal{H}_{\theta}^{\mathrm{ctrl}}(t,x,y,\pi;\,\psi_{x},\psi_{xx},\psi_{xy}):=x\,\pi^{\top}b(y,\theta)\,\psi_{x}+\frac{1}{2}x^{2}\,\pi^{\top}\Sigma(y,\theta)\,\pi\,\psi_{xx}+x\,\pi^{\top}\Sigma_{SY}(y,\theta)\,\psi_{xy}.$ (11)
In the full-information HJB interpretation, we evaluate (11) at $(\psi_{x},\psi_{xx},\psi_{xy})=(V_{x},V_{xx},V_{xy})$. In the Pontryagin interpretation used below, the same affine map is evaluated at the corresponding $\theta$-conditional adjoint/sensitivity objects (denoted $(p^{\theta},p_{x}^{\theta},p_{y}^{\theta})$ along a trajectory); in smooth Markov regimes these objects coincide with the indicated value derivatives.
@@
To see why ex–ante aggregation enters the first-order condition, take any $\theta$-blind admissible perturbation $h=\{h_{t}\}_{t\in[0,T]}$ that is progressively measurable with respect to the observation filtration $(\mathcal{F}_{t}^{\mathrm{obs}})$ and square-integrable, and define $\pi^{\varepsilon}:=\pi+\varepsilon h$ for small $\varepsilon$. For each fixed $\theta$, define $J^{\theta}(\pi):=\mathbb{E}[U(X_{T}^{\pi})\mid\theta]$. The stochastic maximum principle yields the first-variation identity
@@
where $\big(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta}\big)$ denotes the $\theta$-conditional Pontryagin sensitivity/adjoint objects associated with the fixed policy $\pi$ in the frozen-$\theta$ market. (Equivalently, (14) evaluates the same affine map $\partial_{\pi}\mathcal{H}_{\theta}^{\mathrm{ctrl}}$ defined by (11), with $(V_{x},V_{xx},V_{xy})$ replaced by the corresponding Pontryagin objects along the controlled trajectory.)
Frozen-$\theta$ adjoint BSDE and the blocks used in projection.
To make the $\theta$-conditional Pontryagin objects in (14) concrete, we briefly record the backward (stochastic) equation they satisfy for a fixed policy in a frozen-$\theta$ market. Fix a $\theta$-blind feedback policy $\pi\in\mathcal{A}^{\mathrm{fb}}$ and condition on a frozen parameter $\theta$. Let $\mathcal{G}_{t}^{\theta}:=\sigma\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)$ denote the simulator filtration (with the usual augmentation). Under standard hypotheses ensuring well-posedness of the stochastic maximum principle (summarized in Assumption 1), there exists an adapted triple $(p^{\theta},z^{\theta},\tilde{z}^{\theta})$ solving the $\theta$-conditional adjoint BSDE
$p_{t}^{\theta}=U^{\prime}\!\big(X_{T}^{\pi,\theta}\big)+\int_{t}^{T}f_{\theta}\!\big(s,X_{s}^{\pi,\theta},Y_{s}^{\theta},\pi_{s};\,p_{s}^{\theta},z_{s}^{\theta},\tilde{z}_{s}^{\theta}\big)\,ds-\int_{t}^{T}z_{s}^{\theta}\cdot dW_{s}-\int_{t}^{T}\tilde{z}_{s}^{\theta}\cdot dW_{s}^{Y},$ (15)
for a driver $f_{\theta}$ determined by the frozen-$\theta$ dynamics and the fixed policy $\pi$ (e.g. Yong and Zhou, 1999; Pham, 2009). In a smooth Markov regime one may identify $p_{t}^{\theta}=V_{x}^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$ along the controlled state and interpret the blocks entering the portfolio Hamiltonian gradient (11) as
$p_{x,t}^{\theta}:=V_{xx}^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta}),\hskip 18.49988ptp_{y,t}^{\theta}:=V_{xy}^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta}).$ (16)
Equivalently, one may work directly with the second-order stochastic maximum principle, where $(p_{x,t}^{\theta},p_{y,t}^{\theta})$ coincide with the relevant blocks of the second adjoint.
Assumption 1 (Frozen-$\theta$ Markov/BSDE regime (used in Theorems 1 and 2)).
Fix a compact parameter set $\Theta_{0}\subset\Theta$ and a working domain $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$. For every $\theta\in\Theta_{0}$ and every $\theta$-blind feedback policy $\pi$ considered in the paper, assume:
  1. (A1)
    Forward well-posedness and moments. The frozen-$\theta$ forward system (8)–(2) admits a unique strong solution under $\pi$, and for some $p\geq 2$,
    $\sup_{\theta\in\Theta_{0}}\mathbb{E}\Big[\sup_{t\in[0,T]}\big(|X_{t}^{\pi,\theta}|^{p}+\|Y_{t}^{\theta}\|^{p}\big)\Big]<\infty.$
  2. (A2)
    Positivity and utility integrability. Wealth stays strictly positive a.s., and $U$ is $C^{3}$ on $(0,\infty)$ with
    $\sup_{\theta\in\Theta_{0}}\mathbb{E}\Big[|U^{\prime}(X_{T}^{\pi,\theta})|^{2}+|U^{\prime\prime}(X_{T}^{\pi,\theta})|^{2}\Big]<\infty.$
    (For CRRA, this entails suitable negative-moment bounds for $X_{T}^{\pi,\theta}$.)
  3. (A3)
    Adjoint BSDE and smooth blocks. For each $\theta\in\Theta_{0}$, the adjoint BSDE (15) is well posed in $L^{2}$ and admits a Markov decoupling field $u^{\theta}$ on $\mathcal{D}$ of class $C^{1,2}$ in $(t,x,y)$ such that $p_{t}^{\theta}=u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, $p_{x,t}^{\theta}=\partial_{x}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, and $p_{y,t}^{\theta}=\partial_{y}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, with square-integrable bounds.
  4. (A4)
    Nondegenerate driving noise for conditional projections. The block covariance matrix $\Gamma:=\begin{pmatrix}I_{d_{W}}&\rho\\ \rho^{\top}&I_{d_{Y}}\end{pmatrix}$ is positive definite (equivalently, the conditional $L^{2}$ projections used below are taken onto a nondegenerate basis of the driving increments).
  5. (A5)
    Uniformity on compacts. All Lipschitz/growth/nondegeneracy/smoothness constants in (A1)–(A4) can be chosen uniformly for $\theta\in\Theta_{0}$.
Remark 2 (What we assume vs. what we prove).
Assumption 1 collects the analytic conditions required to define the Pontryagin/BSDE objects and to justify the interchanges of limits and conditional expectations used later (notably in the BPTT–BSDE correspondence of Theorem 2). Establishing such smooth decoupling fields and uniform bounds for general high-dimensional coefficients is nontrivial; we do not attempt to do so here. Instead, our contributions are conditional: given a regime where the frozen-$\theta$ SMP/BSDE objects are well behaved (uniformly on compacts), we show that BPTT recovers them in the small-step limit and that the resulting projection yields quantitative residual-based policy-gap control.
Because both $\pi$ and $h$ are $\theta$-blind, taking the outer expectation over $\theta\sim q$ and using Fubini’s theorem gives
@@
Consider the fixed-$q$ ex–ante objective (9) over the $\theta$-blind Markov feedback class $\mathcal{A}^{\mathrm{fb}}$. Assume Assumption 1 holds (in particular, the first-variation identity (14) is valid within $\mathcal{A}^{\mathrm{fb}}$ and the associated $\theta$-conditional Pontryagin/BSDE objects exist). If $\pi^{\star,\mathrm{blind}}$ is a locally optimal interior policy in $\mathcal{A}^{\mathrm{fb}}$, then (18) holds. Moreover, in the portfolio setting (11), the aggregated stationarity is equivalent to the statewise linear system
@@
Fix an interior locally optimal $\theta$-blind feedback $\pi^{\star,\mathrm{blind}}\in\mathcal{A}^{\mathrm{fb}}$. For any $\theta$-blind admissible perturbation $h=\{h_{t}\}_{t\in[0,T]}$ (progressively measurable w.r.t. $(\mathcal{F}_{t}^{\mathrm{obs}})$ and square-integrable), set $\pi^{\varepsilon}:=\pi^{\star,\mathrm{blind}}+\varepsilon h$.
  1. 1.
    First variation and $q$-aggregation. For each fixed $\theta$, the $\theta$-conditional first-variation identity (14) is standard for diffusion control under a frozen parameter (e.g. Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009). Taking the outer expectation over $\theta\sim q$ and using Fubini/Tonelli under the standing integrability assumptions yields (17). Remark. This step implements the fixed-$q$ decision-time commitment model: $\theta$ is sampled only inside the simulator to form the ex–ante objective $J(\cdot)$, while admissible perturbations remain $\theta$-blind. We do not formulate the belief-state/partial-information PMP, where stationarity would be expressed in terms of time-varying posteriors $q_{t}$ (cf. Remark 3).
  2. 2.
    Vanishing first variation $\Rightarrow$ aggregated stationarity. Since $\pi^{\star,\mathrm{blind}}$ is an (interior) local maximizer, the first variation must vanish:
    $\left.\frac{d}{d\varepsilon}J(\pi^{\varepsilon})\right|_{\varepsilon=0}=0\hskip 18.49988pt\text{for all such $\theta$-blind perturbations $h$.}$
    Using standard spike-variation/density arguments within the $\theta$-blind admissible class (i.e., choosing $h_{t}$ supported on arbitrarily small time intervals with arbitrary directions), we obtain the aggregated Pontryagin condition (18), a.s. for a.e. $t\in[0,T]$.
  3. 3.
    Affine-in-control Hamiltonian $\Rightarrow$ linear system and projected form. For the quadratic portfolio Hamiltonian (11), $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$. Substituting the explicit expression for $\partial_{\pi}\mathcal{H}^{\mathrm{ctrl}}_{\theta}$ into (18) and introducing the projection inputs (19)–(21) (and using $x>0$ to divide out a common factor) yields the statewise linear system (22). Whenever $A_{t}(t,x,y)$ is invertible on the working domain, this is equivalently expressed as the projected rule $\pi^{\mathrm{agg}}(t,x,y)=-A_{t}(t,x,y)^{-1}g_{t}(t,x,y)$, i.e. (23).